# MAViS: Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Arrays

### v1

@@ III.4 Barrier Virtualization in the ON-coupling regime
In order to find the non-linear corrections, we use the global tracking method presented in Sec. II.5. We note that, in general, this method is more sensitive to errors in the interdot-identification process, such as misidentified interdots, false positives, and boundary effects, as it does not include a simple way of excluding outliers. However, because we have already performed virtualization in the OFF regime, the large-scale shifts have already been accounted for, which makes boundary effects much less problematic. For instance, the total shift in Fig. 4(a) is much smaller than the corresponding shift in 3(a), despite the barrier being sampled over a much larger range, at about $5$ $\mathrm{mV}$ in the ON regime compared to around $20$ $\mathrm{mV}$ for the OFF regime.
As the barrier is pulsed and the QDs are brought close together, the center of the honeycomb shifts. Particularly, as shown in Fig. 4(c) and Fig. 4(e), the center of the (N5, N8) 2D CSD is seen to follow a quadratic trajectory as barrier J8 is lowered. We observe the same trend across all double QDs. The quadratic and linear compensation coefficients, depicted in Fig. 4(f) and Fig. 4(g), respectively, are derived as described in Sec. II.1. We observe that corrections to the barrier matrix defined in the ON-coupling regime can be of up to $\sim 10$ in the first order. Finally, we note that the square root of the quadratic and linear coefficients are of the same order, and both terms are equally important to virtualizing the barriers.
Having obtained the best-fit coefficients, we then perform barrier pulses that incorporate quadratic and linear compensations on the plungers, by constructing virtualized barriers, Ki for $i\in[1,\dots,12]$, according to equation 5. As the virtual gate Ki is varied, we observe that the map is maintained at the charge symmetry point, allowing for a wide range of barrier pulses, as shown in Fig. 4(b). The compensation matrix values for the linear and quadratic terms, together with measurements on additional double quantum dots, are provided in the Supplemental Material [60].
@@ IV Summary and outlook
Despite significant progress in operating and automating the control of QD devices, gate virtualization remains a challenging and time-consuming process that hinders high-level control of spin qubit arrays. In this work, we have introduced MAViS, a modular and scalable framework that enables the autonomous construction of a complete stack of multi-layer virtual gates starting from raw plunger and barrier gate voltages and sensor readouts in arbitrary architectures. We benchmark MAViS on a 10 quantum dot array with 2D connectivity and demonstrate that full virtualization can be achieved in only $\approx 5$ $\mathrm{h}$, inclusive of both data acquisition and processing (see Supplemental Material [60]).
Motivated by the need for precise control over key Hamiltonian parameters in spin qubit arrays, our approach addresses the virtualization challenge by applying modern ML techniques to state-of-the-art QD arrays. We envision our approach to be useful also by other platforms such as in Kitaev chains [61, 62, 63], in Andreev qubits [64, 65] and in topological readout schemes [66, 67].
@@
Likewise, we highlight the generalizability of the ML models. Despite being trained using exclusively simulated, unvirtualized data [35], the model ensemble is able to decode features from charge stability diagrams both before and after the plunger orthogonalization and normalization stages. Additionally, the models are able to correctly classify CSDs from a 2D array of germanium hole qubits, even though its training data was simulated for electrons in a 1D quantum nanowire, thus demonstrating that our tools are device-agnostic. Moreover, it is important to highlight that the development and deployment of tools based on ML and signal processing (Hough transform, regression models, etc.) has allowed us to track interdots and map the full dependence of the charge states with barrier voltages. Furthermore, numerical simulations revealed that the observed beyond-linear dependence can be caused by an increase in capacitive coupling between the barrier gate and the QDs.
@@
While our virtualization flow has succeeded in virtualizing a large array, we note that several potential points of failure remain. First, it is possible that the CSD itself might be too noisy for the ML model, especially for QDs near the center of the device, which can be more difficult for the sensors to pick up. In our case, because the device has multiple sensors, we found that when one sensor produced a sub-optimal CSD, one of the other sensors produced a CSD that was clearer with distinct transitions. By running the analysis on the output from all four sensors and selecting the clearest result, we were able to handle cases where some of the sensors gave noisy data. Secondly, the pixel classifier occasionally misidentified transitions or identified false positives, particularly in cases with large latching effects. This also occurred where there was a large gradient across the region defined by a specific charge state, such as the central hexagon in Fig. 2(a), where the pixel-classifier marked a red line in the center of the image which does not correspond to any transition. In both cases, we mitigated these errors by carefully designing the postprocessing methods to be robust to such errors.
Figure 5: Simulated barrier virtualization in the ON regime, showing the widening of the interdot region and the quadratic shift in the charge state center. (a, b) Simulated charge stability diagrams illustrating the shift in interdot transitions (cyan dots) and the center of the $(1,1)$ charge state (magenta diamonds) as a function of the virtualized barrier in the OFF ($B$) and ON ($J$) regimes, respectively. (c, d) Overlaid positions of the interdot transitions (cyan) and the [1, 1] charge state center (purple) as functions of $J$ and $K$. (e) The quadratic fit to the position of the $(1,1)$ charge state center as a function of $J$ and $K$.
@@
While MAViS enables the virtualization of the barrier gates with respect to the QD charge states, an end-to-end virtualization stack would require compensating for the effect of each barrier to every interdot tunnel coupling. We note that this effect can be studied with precision by mapping the exchange coupling as measured via resonant qubit spectroscopy as a function of virtualized barriers in the ON regime [68, 7]. This aspect does not require decoding further information from charge stability maps, which has been the central focus of our work.
@@
The extracted capacitive coupling matrices obtained from our methods contain rich information about the device and can be analyzed further to gain insight into the effective location of the QDs, the disorder landscape, and the general impurity density over each metallic gate. This approach will allow the community to autonomously track quantum-dot features for calibration of large-scale quantum-dot arrays.
@@ IV.1 Data Availability Statement
The experimental datasets acquired and analyzed during the current study are available via Zenodo [69]. The processed data, as well as complete figure source files, will be made publicly available upon publication.
Acknowledgements.
The authors are grateful to S. de Snoo for support in the control software and to M. Gullans, C. White, S. Gandhari, and all the members of the Veldhorst lab for fruitful discussions. This research was sponsored in part by the Army Research Office (ARO) under Awards No. W911NF-23-1-0110 and W911NF-23-1-0258. We acknowledge support from the European Union through the IGNITE project with grant agreement No. 101069515 and from the Dutch Research Council (NWO) via the National Growth Fund program Quantum Delta NL (Grant No. NGF.1582.22.001). The views, conclusions, and recommendations contained in this paper are those of the authors and are not necessarily endorsed nor should they be interpreted as representing the official policies, either expressed or implied, of the Army Research Office (ARO) or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by the National Institute of Standards and Technology.
@@ Appendix A Capacitative simulations of quadratic behavior
The quadratic shift observed in the interdot position with decreasing barrier voltage, as presented in Fig. 4, can be reproduced using a capacitive model, where the capacitive couplings are given gate voltage dependence. In particular, we expect that more negative barrier voltages should draw the dots together, leading to an increased dot-dot and barrier-dot capacitive coupling. Our simulations show that the increase in barrier voltage as the barrier gate voltage decreases is sufficient to account for the quadratic dependence we observed, whilst the increase in the dot-dot coupling widens the interdot.
@@
Fig. 5 shows a simulated recreation of the barrier virtualization into the ON regime, qualitatively reproducing the observed behavior in Fig. 4. These simulations were performed using the open-source package QArray [70]. The system is modeled as a double QD controlled by two plunger gates ($\mathrm{P}_{1}$, $\mathrm{P}_{2}$) and one barrier gate ($\mathrm{B}$). The capacitive couplings are captured by the following capacitance matrices:
$C_{dd}(\mathrm{B})$ $=\begin{bmatrix}0&0.5b\\ 0.5b&0\end{bmatrix}$ (8)
$C_{gd}(\mathrm{B})$ $=0.05\begin{bmatrix}1&0&1+0.8b\\ 0&1&1\end{bmatrix}$ (9)
@@
where $C_{dd}$ and $C_{gd}$ represent the dot-dot and gate-dot capacitive couplings, respectively, which we allow to depend upon the barrier gate voltage through, $b=-\mathrm{B}/10^{3}$. We construct the normalized plunger gate voltages, $\mathrm{N}_{1}$ and $\mathrm{N}_{2}$, at $\mathrm{B}=0$ $\mathrm{mV}$ and define the OFF regime virtualized barrier $\mathrm{J}$ at $\mathrm{B}=-25$ $\mathrm{mV}$. In the ON regime, the virtualized barrier $\mathrm{K}$ is defined by fitting a quadratic curve to the position of the charge state center as a function of the barrier voltage. We find that linear trends, such as shown in Fig. 4(e), are indicative of errors in the linear virtualization coefficients. For qualitative agreement with Fig. 4(e), we include an error in the barrier virtualization against $\mathrm{N}_{2}$, so that the coefficient was taken as $-0.96$ rather than the optimal value of $1.00$. This model provides a clear capacitative interpretation of the quadratic trends observed experimentally.
Supplemental Material: MAViS: Modular Autonomous Virtualization System
for Two-Dimensional Semiconductor Quantum Dot Arrays
Appendix Supplementary Section 1 MAViS: scalability and complexity
Our proposed automated virtualization stack (MAViS) is general and applicable to arbitrary quantum dot (QD) device architectures. At its core, MAViS derives the virtual plunger and virtual barrier gates by processing two-dimensional (2D) charge stability diagrams (CSDs) (referred to as “maps” hereon) measured in plunger gates space with the barrier gates adjusted as necessary. Here, we present the complexity analysis of the MAViS framework. We discuss how the number of maps required to virtualize a device scales with the number of metallic gates for one-dimensional (1D) QD arrays before shifting focus to non-1D cases.
Supplementary Section 1.1 One-dimensional QD arrays
Consider a QD device with $N_{\mathrm{S}}$ charge-sensing plunger gates $\mathbf{S}$, $N_{\mathrm{P}}$ plunger gates $\mathbf{P}$, and $N_{\mathrm{B}}$ barrier gates $\mathbf{B}$. The first layer, charge sensor virtualization, in MAViS requires $(N_{\mathrm{S}}\times N_{\mathrm{P}}+N_{\mathrm{S}}\times N_{\mathrm{B}})$ maps. Since $N_{\mathrm{B}}\approx N_{\mathrm{P}}+1$ in most linear arrays, virtualizing sensors requires $O(N_{\mathrm{P}}\,N_{\mathrm{S}})$ maps. However, since the magnitude of compensations decays with the distance between the gates, we can avoid measuring gate pairs located farther than a maximum distance. Considering an $n_{S}$-nearest neighbor approach, where $n_{S}\ll N_{\mathrm{P}}$ and only the $n=\{1,2,\dots n_{S}\}$ nearest neighboring metallic gates in a 1D QD array share a cross-capacitance, the capacitance vanishes with the $(n_{S}+1)$ neighbor and we can reduce the number of required maps to $O(n_{S}\,N_{\mathrm{S}})\approx O(N_{\mathrm{S}})$.
When orthogonalizing plungers (layer 2 in MAViS), a single map suffices to extract the capacitive coupling between two neighboring plungers, which correspond to two coefficients in the $(N_{\mathrm{P}}\times N_{\mathrm{P}})$ plunger-plunger compensation matrix. Thus, the all-to-all plunger connectivity requires $N_{\mathrm{P}}(N_{\mathrm{P}}-1)/2=O(N_{\mathrm{P}}^{2})$ maps. However, following the same $n_{P}$-nearest neighbor approach, the total number of maps for plunger-orthogonalization is $O(N_{\mathrm{P}})$. Next, to perform normalization (layer 3) we would again require $O(N_{\mathrm{P}})$ maps.
Barrier virtualization (steps 4 and 5 in MAViS) involves tracking the center of a selected honeycomb in a series of 2D CSDs and fitting a curve to the resulting trajectory. Assuming that each barrier is stepped $m$ times during barrier virtualization, the all-to-all coupling approach would necessitate $[N_{\mathrm{P}}(N_{\mathrm{P}}-1)\times N_{\mathrm{P}}\times m]/2$ maps. However, since the capacitive coupling of barriers is restricted to the nearest plungers, we only need to worry about the number of (nearest) neighbors of the barriers, $n_{B}$, that affect the crosstalk while others can be set to zero. In this case, the number of maps required is $[n_{B}(n_{B}-1)N_{\mathrm{P}}m]=O(m\,N_{\mathrm{P}})$ both in the low- and high-tunnel coupling regime.
Combining all the steps of MAViS, we obtain a linear scaling for virtualizing sensors and plungers with a constant prefactor given by the number of neighbors with non-vanishing cross-capacitances. For virtualizing barrier gates, we would again require $O(N_{\mathrm{P}})$ maps with the prefactor given by the number of barrier steps that depend on the experiment and the number of nearest neighbors. Thus, the virtualization framework scales linearly with the number of plungers and barriers for a 1D device.
Supplementary Section 1.2 Higher-dimensional QD arrays
MAViS would follow the same scaling in higher dimensions, however, the number of nearest neighbors $\{n_{S},n_{P},n_{B}\}$ will be dimension dependent. Thus, performing MAViS on a higher dimensional QD device would require $O(n_{S}\times~{}s)$, $O(n_{P}\times N_{\mathrm{P}}\times m)$, $O(n_{B}\times N_{\mathrm{B}}\times m)$ maps for virtualization of sensors, plungers, and barriers, respectively.
In the case of the ten-QD two-dimensional device used in this work, the estimated average time necessary to acquire and process the experimental data with MAViS, as well as the total number of measurements for each virtualization layer, are detailed in Table 1. In layer 2, we have only considered the cross capacitance compensations up to the second nearest neighbors. We neglected cross capacitances between elements that are very far in the array design. For instance, we assume that the cross capacitance of $\mathrm{P}_{4}$ to $\mathrm{P}_{2}$ and $\mathrm{P}_{4}$ to $\mathrm{P}_{9}$ is negligible. Furthermore, we note that measurements in the barrier fine virtualization step are highly oversampled, with maps acquired for every $0.5$ $\mathrm{mV}$ shift in the barrier voltage. This was done to enable a highly precise fitting in the high-coupling regime. In practice, $O(10)$ samples are sufficient to characterize the non-linear trend and only maps between neighboring plungers and barriers suffice for barrier virtualization in the low-coupling regime. Taking all of this into account, the total time MAViS would require to virtualize the ten-dot device would be about $2$ $\mathrm{h}$ $15$ $\mathrm{min}$ ($8,064$ $\mathrm{s}$) as opposed to $5$ $\mathrm{h}$ $17$ $\mathrm{min}$ ($19,020$ $\mathrm{s}$) suggested by Table 1.
Supplementary Table 1: Modular Automated Virtualization System (MAViS) timing information when applied to the ten-QD device. The estimated mean data acquisition and MAViS processing times are given per map and are calculated based on a total measurement and processing, respectively, time divided by the number of maps in a given layer. For the barrier fine virtualization (layer 5), a large number of maps was acquired to carefully follow the non-linear dependence. However, in practice, only a small fraction of the total number of measurements is required.
Virt. layer MAViS step Data acquisition time [s] MAViS processing time [s] Total no. of measurements
1 Charge sensor virtualization 0.34 0.002 88
2 Plunger gate orthogonalization 2.3 35.3 27
3 Plunger gate normalization 2.4 35.3 10
4 Barrier coarse virtualization 4.5 3.9 672
5 Barrier fine virtualization 4.5 3.8 1440
Appendix Supplementary Section 2 Additional figures: Detailed capacitive coefficients in the off-regime
Supplementary Figure 1: Charge sensor virtualization. Section of $M_{1}^{-1}$ characterizing the relative cross-capacitance of each gate to the four charge-sensing plunger gates. The full $M_{1}^{-1}$ is not shown, as, in practice, the full matrix is a $26\times 26$ matrix with diagonal entries set to 1, similar to what is shown in the Supplementary Information of Refs. [20,25].
Supplementary Figure 2: (a) Plunger gate orthogonalization as depicted in Fig. 2(d). Compensation matrix $M_{2}^{-1}$ obtained from MAViS for plunger orthogonalization. (b) Plunger gate normalization depicted in Fig. 2(f). Compensation matrix $M_{3}^{-1}$ obtained from MAViS for plunger normalization.
Supplementary Figure 3: (a) Section of $M_{4}^{-1}$ showing the barrier-to-plunger virtualization coefficients in the off-coupling regime as depicted in Fig. 3(g). (b) Section of $M_{4}^{-1}$ including refinements obtained from additional three rounds of measurements. The effect of additional rounds of correction is, on average, $-0.07(8)$ for the first round, $-0.002(36)$ for the second round, and $-0.007(34)$ for the third round. These numbers represent the mean and standard deviation of the full corrections matrices, and their converging trend to a correction matrix with mean of $\approx 0$ suggests two rounds of correction are sufficient for a reliable barrier virtualization. The standard deviation of the last two correction matrices ($\approx 0.03$) can be viewed as an average detection error of our algorithm.
Appendix Supplementary Section 3 Additional figures: Detailed capacitive coefficients in the on-regime
Supplementary Figure 4: Barrier virtualization in the on-coupling regime depicted in Fig. 4(f). (a) Matrix showing the square root of the quadratic coefficients ($\sqrt{\alpha}$) in the on-coupling regime. (b) Matrix illustrating the linear coefficients ($\beta$) in the on-coupling regime.
Appendix Supplementary Section 4 Additional figures: Barrier virtualization in the on-coupling regime
Supplementary Figure 5: Virtualizing gate ${\mathrm{J}}_{2}$. (a) Sequence of CSDs N5 vs N1 as a function of J2 in the range [5, -85] $\mathrm{mV}$ with respect to the dc reference point. (b) Sequence of CSDs N5 vs N1 as a function of K2 in the same range. The cyan points in (a) and (b) indicate the position of the interdots returned by the ML module and the magenta diamonds indicate the center of the tracked honeycomb. The shift in the (N1, N5) space of the central charge sector in panel (a) reveals imperfect virtualization over the large voltage range (90 $\mathrm{mV}$). Panel (b), where the finely virtualized gate K2 is adopted, shows a decreased susceptibility to the barrier gate voltage change. (c, d) Concise presentation of the evolution of the charge state while stepping J2 and K2, respectively. The finely calibrated K2 preserves the position of the charge state. (e) The fit to the center of the honeycomb positions for plunger gates N1 and N5 as a function of barrier gate J2 indicating a beyond-linear dependence for N1. (f, g) The quadratic coefficient $\sqrt{\alpha}$ and linear compensation coefficient $\beta$ respectively, with K2 vs. N1 and vs. N5 highlighted.
Supplementary Figure 6: Virtualizing gate ${\mathrm{J}}_{5}$. (a) Sequence of CSDs N6 vs N3 as a function of J5 in the range [5, -55] $\mathrm{mV}$ with respect to the dc reference point. (b) Sequence of CSDs N6 vs N3 as a function of K5 in the same range. The cyan points in (a) and (b) indicate the position of the interdots returned by the ML module and the magenta diamonds indicate the center of the tracked honeycomb. The shift in the (N3, N6) space of the central charge sector in panel (a) reveals imperfect virtualization over the large voltage range (60 $\mathrm{mV}$). Panel (b), where the finely virtualized gate K5 is adopted, shows a decreased susceptibility to the barrier gate voltage change. (c, d) Concise presentation of the evolution of the charge state while stepping J5 and K5, respectively. The finely calibrated K5 preserves the position of the charge state. (e) The fit to the center of the honeycomb positions for plunger gates N3 and N6 as a function of barrier gate J5 indicating a beyond-linear dependence for N3. (f, g) The quadratic coefficient $\sqrt{\alpha}$ and linear compensation coefficient $\beta$ respectively, with K5 vs. N3 and vs. N6 highlighted.
Supplementary Figure 7: Virtualizing gate ${\mathrm{J}}_{10}$ (a) Sequence of CSDs N6 vs N9 as a function of J10 in the range [20, -25] $\mathrm{mV}$ with respect to the dc reference point. (b) Sequence of CSDs N6 vs N9 as a function of K10 in the same range. The cyan points in (a) and (b) indicate the position of the interdots returned by the ML module and the magenta diamonds indicate the center of the tracked honeycomb. The shift in the (N9, N6) space of the central charge sector in panel (a) reveals imperfect virtualization over the large voltage range (60 $\mathrm{mV}$). Panel (b), where the finely virtualized gate K10 is adopted, shows a decreased susceptibility to the barrier gate voltage change. (c, d) Concise presentation of the evolution of the charge state while stepping J10 and K10, respectively. The finely calibrated K10 preserves the position of the charge state. (e) The fit to the center of the honeycomb positions for plunger gates N6 and N9 as a function of barrier gate J10 indicating a beyond-linear dependence for N9. (f, g) The quadratic coefficient $\sqrt{\alpha}$ and linear compensation coefficient $\beta$ respectively, with K10 vs. N6 and vs. N9 highlighted.

### v2

@@ III.4 Barrier virtualization in the ON-coupling regime
In order to find the nonlinear corrections, we use the global tracking method presented in Sec. II.5. We note that, in general, this method is more sensitive to errors in the interdot-identification process, such as misidentified interdots, false positives, and boundary effects, as it does not include a simple way of excluding outliers. However, because we have already performed virtualization in the OFF regime, the large-scale shifts have already been accounted for, which makes boundary effects much less problematic. For instance, the total shift in Fig. 4(a) is much smaller than the corresponding shift in Fig. 3(a), despite the barrier being sampled over a much larger range, at about $5$ $\mathrm{mV}$ in the ON regime compared to around $20$ $\mathrm{mV}$ for the OFF regime.
As the barrier is pulsed and the QDs are brought close together, the center of the honeycomb shifts. Particularly, as shown in Figs. 4(c) and 4(e), the center of the (N5, N8) 2D CSD is seen to follow a quadratic trajectory as barrier J8 is lowered. We observe the same trend across all double QDs. The quadratic and linear compensation coefficients, depicted in Figs. 4(f) and 4(g), respectively, are derived as described in Sec. II.1. We observe that corrections to the barrier matrix defined in the ON-coupling regime can be of up to $\sim 10$ in the first order. Finally, we note that the square root of the quadratic and linear coefficients are of the same order, and both terms are equally important to virtualizing the barriers.
Having obtained the best-fit coefficients, we then perform barrier pulses that incorporate quadratic and linear compensations on the plungers, by constructing virtualized barriers, Ki for $i\in[1,\dots,12]$, according to Eq. (5). As the virtual gate Ki is varied, we observe that the map is maintained at the charge symmetry point, allowing for a wide range of barrier pulses, as shown in Fig. 4(b). The compensation matrix values for the linear and quadratic terms, together with measurements on additional double quantum dots, are provided in Supplemental Material [60].
@@ IV Summary and outlook
Despite significant progress in operating and automating the control of QD devices, gate virtualization remains a challenging and time-consuming process that hinders high-level control of spin qubit arrays. In this work, we have introduced MAViS, a modular and scalable framework that enables the autonomous construction of a complete stack of multilayer virtual gates starting from raw plunger and barrier gate voltages and sensor readouts in arbitrary architectures. We benchmark MAViS on a 10-QD array with 2D connectivity and demonstrate that full virtualization can be achieved in only $\approx 5$ $\mathrm{h}$, inclusive of both data acquisition and processing (see Supplemental Material [60]).
Motivated by the need for precise control over key Hamiltonian parameters in spin qubit arrays, our approach addresses the virtualization challenge by applying modern ML techniques to state-of-the-art QD arrays. We envision our approach to be useful also by other platforms such as in Kitaev chains [61, 62, 63], in Andreev qubits [64, 65], and in topological readout schemes [66, 67].
@@
Likewise, we highlight the generalizability of the ML models. Despite being trained using exclusively simulated, unvirtualized data [35], the model ensemble is able to decode features from charge stability diagrams both before and after the plunger orthogonalization and normalization stages. Additionally, the models are able to correctly classify CSDs from a 2D array of germanium hole qubits, even though its training data were simulated for electrons in a 1D quantum nanowire, thus demonstrating that our tools are device agnostic. Moreover, it is important to highlight that the development and deployment of tools based on ML and signal processing (Hough transform, regression models, etc.) has allowed us to track interdots and map the full dependence of the charge states with barrier voltages. Furthermore, numerical simulations revealed that the observed beyond-linear dependence can be caused by an increase in capacitive coupling between the barrier gate and the QDs.
@@
While our virtualization flow has succeeded in virtualizing a large array, we note that several potential points of failure remain. First, it is possible that the CSD itself might be too noisy for the ML model, especially for QDs near the center of the device, which can be more difficult for the sensors to pick up. In our case, because the device has multiple sensors, we were always able to find a clear CSD with distinct charge transitions. By running the analysis on the output from all four sensors and selecting the clearest result, we were able to handle cases where some of the sensors gave suboptimal data. Second, the pixel classifier occasionally misidentified transitions or identified false positives, particularly in cases with large latching effects. This also occurred where there was a large gradient across the region defined by a specific charge state, such as the central hexagon in Fig. 2(a), where the pixel classifier marked a red line in the center of the image which does not correspond to any transition. In both cases, we mitigated these errors by carefully designing the postprocessing methods to be robust to such errors.
@@
At present, MAViS enables the virtualization of the barrier gates with respect to the QD charge states. An end-to-end virtualization stack would require compensating for the effect of each barrier to every interdot exchange coupling. This effect can be studied with precision by mapping the exchange coupling as measured via resonant qubit spectroscopy as a function of virtualized barriers in the ON regime [68, 7, 51]. Since compensating for exchange coupling requires qubit characterization rather than decoding information from CSDs, it is left for future work.
@@
The extracted capacitive coupling matrices obtained from our methods contain rich information about the device and can be analyzed further to gain insight into the effective location of the QDs, the disorder landscape, and the general impurity density over each metallic gate. This approach will allow the community to autonomously track QD features for calibration of large-scale QD arrays.
Acknowledgements.
The authors are grateful to S. de Snoo for support in the control software and to M. Gullans, C. White, S. Gandhari, and all the members of the Veldhorst lab for fruitful discussions. This research was sponsored in part by the U.S. Army Research Office (ARO) under Awards No. W911NF-23-1-0110 and No. W911NF-23-1-0258. We acknowledge support from the European Union through the IGNITE project with Grant Agreement No. 101069515 and from the Dutch Research Council (NWO) via the National Growth Fund program Quantum Delta NL (Grant No. NGF.1582.22.001). The views, conclusions, and recommendations contained in this paper are those of the authors and are not necessarily endorsed nor should they be interpreted as representing the official policies, either expressed or implied, of the U.S. Army Research Office (ARO) or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by the National Institute of Standards and Technology.
@@ IV.1 Data Availability Statement
The data that support the findings of this article are openly available [69]. Complete figure source files are available via data.nist.gov [70].
@@ Appendix A Capacitative simulations of quadratic behavior
The quadratic shift observed in the interdot position with decreasing barrier voltage, as presented in Fig. 4, can be reproduced using a capacitive model, where the capacitive couplings are given gate voltage dependence. In particular, we expect that more negative barrier voltages should draw the dots together, leading to an increased dot-dot and barrier-dot capacitive coupling. Our simulations show that the increase in barrier voltage as the barrier gate voltage decreases is sufficient to account for the quadratic dependence we observed, while the increase in the dot-dot coupling widens the interdot.
@@
Figure 5 shows a simulated recreation of the barrier virtualization into the ON regime, qualitatively reproducing the observed behavior in Fig. 4. These simulations were performed using the open-source package QArray [71]. The system is modeled as a double QD controlled by two plunger gates ($\mathrm{P}_{1}$, $\mathrm{P}_{2}$) and one barrier gate ($\mathrm{B}$). The capacitive couplings are captured by the following capacitance matrices:
$C_{dd}(\mathrm{B})$ $=\begin{bmatrix}0&0.5b\\ 0.5b&0\end{bmatrix},$ (8)
$C_{gd}(\mathrm{B})$ $=0.05\begin{bmatrix}1&0&1+0.8b\\ 0&1&1\end{bmatrix},$ (9)
@@
where $C_{dd}$ and $C_{gd}$ represent the dot-dot and gate-dot capacitive couplings, respectively, which we allow to depend upon the barrier gate voltage through $b=-\mathrm{B}/10^{3}$. We construct the normalized plunger gate voltages, $\mathrm{N}_{1}$ and $\mathrm{N}_{2}$, at $\mathrm{B}=0$ $\mathrm{mV}$ and define the OFF regime virtualized barrier $\mathrm{J}$ at $\mathrm{B}=-25$ $\mathrm{mV}$. In the ON regime, the virtualized barrier $\mathrm{K}$ is defined by fitting a quadratic curve to the position of the charge state center as a function of the barrier voltage. We find that linear trends, such as shown in Fig. 4(e), are indicative of errors in the linear virtualization coefficients. For qualitative agreement with Fig. 4(e), we include an error in the barrier virtualization against $\mathrm{N}_{2}$, so that the coefficient was taken as $-0.96$ rather than the optimal value of $1.00$. This model provides a clear capacitative interpretation of the quadratic trends observed experimentally.
Figure 5: Simulated barrier virtualization in the ON regime, showing the widening of the interdot region and the quadratic shift in the charge state center. (a), (b) Simulated charge stability diagrams illustrating the shift in interdot transitions (cyan dots) and the center of the $(1,1)$ charge state (magenta diamonds) as a function of the virtualized barrier in the OFF ($\mathrm{B}$) and ON ($\mathrm{J}$) regimes, respectively. (c), (d) Overlaid positions of the interdot transitions (cyan) and the [1, 1] charge state center (purple) as functions of $\mathrm{J}$ and $\mathrm{K}$. (e) The quadratic fit to the position of the $(1,1)$ charge state center as a function of $\mathrm{J}$ and $\mathrm{K}$.